OpenStudy (anonymous):

how do i write a function rule for this table where X is 2, 4, 6 and Y is 1, 0, -1

OpenStudy (amistre64):

the differences look pretty constant, which would suggest a linear equation

OpenStudy (amistre64):

y drops by 1 as x increase by 2 would suggest a slope of -1/2

OpenStudy (amistre64):

y = -1/2 x when y=0, x=4 0 = -4/2 , we need to "add" an adjustment; +2 should fit y = -1/2 x + 2

OpenStudy (anonymous):

thank you very much you are great!!! :)

OpenStudy (amistre64):

lol, good luck